# Solving Football by Exploiting Equilibrium Structure of 2p0s Differential Games with One-Sided Information

### v3

@@ 7 Conclusion
Unlike general IIEFGs where NE strategies are distributions over the entire action space, we showed that 2p0s1 games enjoy a much simpler NE structure when P1 can precisely control the public belief, leading to exponentially reduced game-tree complexity. We demonstrated the utilities of this NE structure in solving games with continuous action spaces in model-free and model-based modes, in terms of computational cost and solution quality. Our methods enable fast approximation of deceptive and counter-deception team strategies, e.g., in sports, missile/drone defense, and risk-sensitive robotics applications, tailored for specific team dynamics, action spaces, and task specifications.
@@ Appendix A Connection between Primal and Dual Games
Here we use the infinitely-repeated game setting to explain the connection between the primal and dual games and the interpretation of the dual variable $\hat{p}$. Please see Theorem 2.2 in De Meyer (1996) and the extension to differential games in Cardaliaguet (2007).
@@
In an infinitely-repeated normal-form game with one-sided information, there is a set of $I$ possible payoff tables $\{A_{i}\}_{i=1}^{I}$. A table is initially drawn from $p_{0}$ and shown to P1, while P2 only knows $p_{0}$. At stage $t\in[T]$, the players draw actions from their strategies $\eta_{i}$ and $\zeta$, respectively, under the current belief $p_{t}$. The actions are public, which trigger belief updates, yet the actual payoffs are hidden (or otherwise P2 may infer the true payoff easily). At the terminal $T$, an average payoff is received by P1 from P2. This game is proven to have a value (Aumann et al., 1995) for finite and infinite $T$. Here we only consider the latter case where the value is defined as
@@
$\min_{\{u^{k}\},\{\alpha_{ki}\}}\max_{\{v^{k}\}}\qquad\sum_{k=1}^{I}\lambda^{k}\Big(V(t+\tau,x^{k},p^{k})+\tau\mathbb{E}_{i\sim p^{k}}[l_{i}(u^{k},v^{k})]\Big)$ ($\text{P}_{1}$)
$\text{s.t.}\qquad u^{k}\in\mathcal{U},\qquad x^{k}=\text{ODE}(x,\tau,u^{k},v^{k};f),\qquad v^{k}\in\mathcal{V},$
$\qquad\qquad\alpha_{ki}\in[0,1],\;\sum_{k=1}^{I}\alpha_{ki}=1,\;\lambda^{k}=\sum_{i=1}^{I}\alpha_{ki}p[i],$
$\qquad\qquad p^{k}[i]=\frac{\alpha_{ki}p[i]}{\lambda^{k}},\qquad\forall i,k\in[I].$
@@
$\min_{\{v^{k}\},\{\lambda^{k}\},\{\hat{p}^{k}\}}\max_{\{u^{k}\}}\sum_{k=1}^{I+1}\lambda^{k}\left(V^{*}(t+\tau,x^{k},\hat{p}^{k}-\tau l(u^{k},v^{k}))\right)$ ($\text{P}_{2}$)
$\text{s.t.}\qquad u^{k}\in\mathcal{U},\qquad v^{k}\in\mathcal{V},\qquad x^{k}=\text{ODE}(x,\tau,u^{k},v^{k};f),$
$\qquad\qquad\lambda^{k}\in[0,1],\;\sum_{k=1}^{I+1}\lambda^{k}\hat{p}^{k}=\hat{p},\;\sum_{k=1}^{I+1}\lambda^{k}=1,\;k\in[I+1].$
@@
$V_{\tau}(t_{0},x_{0},p)$ $=\min_{\{\eta_{i}\}}\mathbb{E}_{i\sim p,~u\sim\eta^{i}}\left[\max_{v\in\mathcal{V}}\left\{V_{\tau}(t_{0}+\tau,x^{\prime}(u,v),p^{\prime}(u))+\tau l_{i}(u,v)\right\}\right]$ (10)
$=\min_{\{\eta_{i}\}}\int_{\mathcal{U}}\bar{\eta}(u)\max_{v\in\mathcal{V}}\left\{V_{\tau}(t_{0}+\tau,x^{\prime}(u,v),p^{\prime}(u))+\tau\mathbb{E}_{i\sim p^{\prime}(u)}[l_{i}(u,v)]\right\}du$
@@
$a(u,p^{\prime}(u))=\max_{v\in\mathcal{V}}V_{\tau}(t_{0}+\tau,x^{\prime}(u,v),p^{\prime}(u))+\tau\mathbb{E}_{i\sim p^{\prime}(u)}[l_{i}(u,v)].$ (11)
@@
For any $(t,x,p)\in[0,T]\times\mathcal{X}\times\Delta(I)$, if A1-5 hold, then there exist $C>0$, such that
$V(t,x,p)\leq\max_{\zeta\in\mathcal{Z}}J(t,x,p;\{\eta_{i,{\tau}}\}^{\dagger},\zeta)\leq V(t,x,p)+C(T-t)\tau.$ (13)
@@
$V^{*}(t,x,\hat{p})\leq\max_{\{\eta_{i}\}\in\{\mathcal{H}^{i}\}^{I}}J^{*}(t,x,\hat{p};\{\eta_{i}\},\zeta_{\tau}^{\dagger})\leq V^{*}(t,x,\hat{p})+C(T-t)\tau.$ (14)
@@ C.1 Settings
where $\tilde{V}_{\tau}(t,x,p^{k})=\min_{u\in\mathcal{U}}\max_{v\in\mathcal{V}}V_{\tau}(t+\tau,x+\tau f(x,u,v),p^{k})+\mathbb{E}_{i\sim p^{k}}[l_{i}(u,v)]$ is the non-revealing value at $(t,x,p^{k})$, and $\sum_{k=1}^{I}\lambda^{k}=1,~\sum_{k=1}^{I}\lambda^{k}p^{k}=p,~p^{k}\in\Delta(I),~\forall k\in[I]$.
From Cardaliaguet (2007), the value $V$ of the original game $G$ is the unique viscosity solution to the following Hamilton-Jacobi equation
@@ C.2 Preliminary lemmas
Lemma C.1 (Value properties).
$V$ is spatially Lipschitz and $\omega$-semiconcave.
@@
Spatial Lipschitzness of $V$ is proved in Cardaliaguet (2007). A function $V:\mathcal{X}\rightarrow\mathbb{R}$ is $\omega$-semiconcave on $\mathcal{X}$ if
$V(y)\leq V(x)+p^{\top}(y-x)+\frac{\omega}{2}\|y-x\|^{2}_{2}\quad\forall x,y\in\mathcal{X},~p\in\partial^{+}V,$ (17)
where $\partial^{+}V$ is the supergradient of $V$. Verbally, between two points on its support, $V$ bends up at most quadratically with a curvature $\omega$. Semiconcavity of value functions for first-order Hamilton–Jacobi equations is proved in Bardi et al. (1997) and Cannarsa & Sinestrari (2004) (Thm. 3.3.7). ∎
Lemma C.2 (Quadratic contact).
For any $(u,v,p)$ with $\|f(x,u,v)\|_{2}\leq F$ and $p\in\Delta(I)$, let $\phi\in C^{2}$ be a test function touching $V$ at $(t,x,p)$ (a local max. or min.). Then
$\bigl|V(t+\tau,x+\tau f,p)-\phi(t+\tau,x+\tau f,p)\bigr|\;\leq\;C\,\tau^{2},\qquad C:=\frac{1}{2}\bigl(\omega+\|\nabla^{2}_{x}\phi\|_{\infty}\bigr)(1+F)^{2}.$ (18)
@@
By semiconcavity of $V$ in $x$ there exists super-derivative $\hat{p}\in\partial^{+}_{x}V(t,x,p)$ with $\hat{p}=\nabla_{x}\phi(t,x,p)$. For $\Delta:=\tau f(x,u,v)$,
$V(t+\tau,x+\Delta,p)-\phi(t+\tau,x+\Delta,p)\leq\frac{\omega}{2}\,\|\Delta\|_{2}^{2}+\frac{1}{2}\|\nabla^{2}_{x}\phi\|_{\infty}(\tau+\|\Delta\|_{2})^{2}.$
The same bound with $V$ and $\phi$ swapped proves Eq. 18. ∎
Lemma C.3 (1‐Lipschitz property).
Let $g,h$ be two bounded functions on $\Delta(I)$. Let $Vex(g)$ be the convex envelope of $g$. Then for any $p\in\Delta(I)$
$|Vex(g)(p)-Vex(h)(p)|\;\leq\;\sup_{p}|g(p)-h(p)|.$
@@
Let $\delta:=\sup_{p}|g(p)-h(p)|$, i.e., $g-\delta\leq h\leq g+\delta$ for all $p$. Then $Vex(g-\delta)\leq Vex(h)\leq Vex(g+\delta)$. Since $Vex(g-\delta)=Vex(g)-\delta$, we have $|Vex(g)-Vex(h)|\leq\delta=\sup_{p}|g(p)-h(p)|$. ∎
Lemma C.4 (Quadratic local error).
For every touching $C^{2}$ test function $\phi$ and every $(t,x,p)\in[0,T-\tau]\times\mathcal{X}\times\Delta(I)$
$\bigl|\,T_{\tau}[V](t,x,p)-T_{\tau}[\phi](t,x,p)\bigr|\;\leq\;C\tau^{2},$
with the constant $C$ independent of $(t,x,p)$.

### v4

@@ 7 Conclusion
Unlike general IIEFGs where NE strategies are distributions over the entire action space, we showed that 2p0s1 games enjoy an atomic NE structure when P1 can precisely control the public belief, leading to an exponentially reduced game-tree complexity. We demonstrated the utilities of this NE structure in solving games with continuous action spaces in model-free and model-based modes, in terms of computational cost and solution quality. Our methods enable fast approximation of deceptive and counter-deception team strategies, with potential applications to sports, missile/drone defense, and risk-sensitive robotics applications, tailored for specific team dynamics, action spaces, and task specifications.
Ethics Statement
LLM Usage.
Large language models (LLMs) were sparingly used for (1) polishing the writing, (2) code generation for visualization, and (3) some proof steps for Thm. 4.1 and Thm. 4.2. All codes and proofs are verified by authors.
Impact.
This work examines strategic behavior in sequential decision-making and how deceptive tactics can arise in autonomous agents. The findings have dual-use implications; all experiments are simulation-only and involve no human subjects or personal data.
Reproducibility Statement
The paper provides all the relevant details needed to setup the experiments and reproduce the results. These artifacts include: 1) value approximation algorithm and the necessary assumptions to achieve the theoretical guarantees, 2) detailed description of the game setups in the appendix, and, 3) fully executable code with adequate documentation shared via a github repository.
@@ Appendix A Connection between Primal and Dual Games
Here we use the infinitely-repeated game setting to explain the connection between the primal and dual games and the interpretation of the dual variable $\hat{p}$. Please see Theorem 2.2 in de1996repeated and the extension to differential games in cardaliaguet2007differential.
@@
In an infinitely-repeated normal-form game with one-sided information, there is a set of $I$ possible payoff tables $\{A_{i}\}_{i=1}^{I}$. A table is initially drawn from $p_{0}$ and shown to P1, while P2 only knows $p_{0}$. At stage $t\in[T]$, the players draw actions from their strategies $\eta_{i}$ and $\zeta$, respectively, under the current belief $p_{t}$. The actions are public, which trigger belief updates, yet the actual payoffs are hidden (or otherwise P2 may infer the true payoff easily). At the terminal $T$, an average payoff is received by P1 from P2. This game is proven to have a value (aumann1995repeated) for finite and infinite $T$. Here we only consider the latter case where the value is defined as
@@
$\min_{\{u^{k}\},\{\alpha_{ki}\}}\max_{\{v^{k}\}}\quad\sum_{k=1}^{I}\lambda^{k}\Big(V(t+\tau,x^{k},p^{k})+\tau\mathbb{E}_{i\sim p^{k}}[l_{i}(x,u^{k},v^{k})]\Big)$ ($\text{P}_{1}$)
$\text{s.t.}\quad u^{k}\in\mathcal{U},\quad x^{k}=x+\tau f(x,u^{k},v^{k}),\quad v^{k}\in\mathcal{V},$
$\quad\quad\alpha_{ki}\in[0,1],\;\sum_{k=1}^{I}\alpha_{ki}=1,\;\lambda^{k}=\sum_{i=1}^{I}\alpha_{ki}p[i],$
$\quad\quad p^{k}[i]=\frac{\alpha_{ki}p[i]}{\lambda^{k}},\quad\forall i,k\in[I].$
@@
$\min_{\{v^{k}\},\{\lambda^{k}\},\{\hat{p}^{k}\}}\max_{\{u^{k}\}}\sum_{k=1}^{I+1}\lambda^{k}\left(V^{*}(t+\tau,x^{k},\hat{p}^{k}-\tau l(x,u^{k},v^{k}))\right)$ ($\text{P}_{2}$)
$\text{s.t.}\quad u^{k}\in\mathcal{U},\quad v^{k}\in\mathcal{V},\quad x^{k}=x+\tau f(x,u^{k},v^{k}),$
$\quad\quad\lambda^{k}\in[0,1],\;\sum_{k=1}^{I+1}\lambda^{k}\hat{p}^{k}=\hat{p},\;\sum_{k=1}^{I+1}\lambda^{k}=1,\;k\in[I+1].$
@@
$V_{\tau}(t_{0},x_{0},p)$ $=\min_{\{\eta_{i}\}}\mathbb{E}_{i\sim p,~u\sim\eta^{i}}\left[\max_{v\in\mathcal{V}}\left\{V_{\tau}(t_{0}+\tau,x^{\prime}(u,v),p^{\prime}(u))+\tau l_{i}(x,u,v)\right\}\right]$ (10)
$=\min_{\{\eta_{i}\}}\int_{\mathcal{U}}\bar{\eta}(u)\max_{v\in\mathcal{V}}\left\{V_{\tau}(t_{0}+\tau,x^{\prime}(u,v),p^{\prime}(u))+\tau\mathbb{E}_{i\sim p^{\prime}(u)}[l_{i}(x,u,v)]\right\}du$
@@
$a(u,p^{\prime}(u))=\max_{v\in\mathcal{V}}V_{\tau}(t_{0}+\tau,x^{\prime}(u,v),p^{\prime}(u))+\tau\mathbb{E}_{i\sim p^{\prime}(u)}[l_{i}(x,u,v)].$ (11)
@@
For any $(t,x,p)\in[0,T]\times\mathcal{X}\times\Delta(I)$, if A1-5 hold, then for any $\epsilon>0$, there exist $C>0$, and $\tau\in(0,\tau^{*}]$, such that
$0\leq\max_{\zeta\in\mathcal{Z}}J(t,x,p;\{\eta_{i,{\tau}}^{\dagger}\},\zeta)-V(t,x,p)\leq\epsilon$ (13)
@@
$0\leq\max_{\{\eta_{i}\}\in\{\mathcal{H}^{i}\}^{I}}J^{*}(t,x,\hat{p};\{\eta_{i}\},\zeta_{\tau}^{\dagger})-V^{*}(t,x,\hat{p})\leq\epsilon$ (14)
@@ C.1 Settings
where $\tilde{V}_{\tau}(t,x,p^{k})=\min_{u\in\mathcal{U}}\max_{v\in\mathcal{V}}V_{\tau}(t+\tau,x+\tau f(x,u,v),p^{k})+\tau\mathbb{E}_{i\sim p^{k}}[l_{i}(x,u,v)]$ is the non-revealing value at $(t,x,p^{k})$, and $\sum_{k=1}^{I}\lambda^{k}=1,~\sum_{k=1}^{I}\lambda^{k}p^{k}=p,~p^{k}\in\Delta(I),~\forall k\in[I]$.
From cardaliaguet2007differential, the value $V$ of the original game $G$ is the unique viscosity solution to the following Hamilton-Jacobi equation
@@ C.2 Preliminary lemmas
Lemma C.1 (Uniform Convergence (Thm 3.1 of cardaliaguet2009numerical).
Assuming (A1-A5) hold, the map $V_{\tau}$ converges uniformly to the value function $V$ on compact subsects of $0,T]\times\mathcal{X}\times\Delta(I)$, in the following sense:
$\lim_{\tau\rightarrow 0^{+},\;t_{k}\rightarrow t,\;x^{\prime}\rightarrow x,\;p^{\prime}\rightarrow p}V_{\tau}(t_{k},x^{\prime},p^{\prime})=V(t,x,p)\quad\forall(t,x,p)\in[0,T]\times\mathcal{X}\times\Delta(I).$
@@
Proved in cardaliaguet2009numerical for both $V_{\tau}$ and $V^{*}_{\tau}$. ∎
Lemma C.2.
Let $\bar{\mathcal{V}}(t_{0})=\{v:[t_{0},T]\rightarrow\mathcal{V}\}$ denote a set of open-loop controls for P2. Then, given $f$, a constant control $u\in\mathcal{U}$ on $t_{k},t_{k+1}]$, a time-varying control $v\in\bar{\mathcal{V}}(t)$, there exists a $w$ such that
$w\in\text{Co}f(x,u,\mathcal{V})$ (17)
where Co denotes the convex hull.
@@
Let
$w:=\frac{1}{\tau}\int_{t_{k}}^{t_{k+1}}f(x,u,v(s))\;ds$ (18)
We show that $w$ as defined above lies in the convex hull of the set $f(x,u,\mathcal{V})$.
Let $S:=f(x,u,\mathcal{V})=\{f(x,u,\bar{v}):\bar{v}\in\mathcal{V}\}$. For any time $s$, since $v(s)\in\mathcal{V}$, we have:
$f(x,u,v(s))\in S\quad\text{ for all }s\in[t_{k},t_{k+1}]$ (19)
Case 1: when $v(\cdot)$ is piecewise-constant.
Assume first that $v(\cdot)$ is piecewise constant on the partition $t_{k},t_{k+1}]$:
$t_{k},t_{k+1}]=\bigcup_{j=1}^{m}N_{j},\quad|N_{j}|=\Delta t_{j},\quad\sum_{j=1}^{m}\Delta t_{j}=\tau,$
and $v(s)=v_{j}$ on $N_{j}$. Then
$\int_{t_{k}}^{t_{k+1}}f(x,u,v(s))ds=\sum_{j=1}^{m}\Delta t_{j}f(x,u,v_{j}).$
Divide by $\tau$ to get
$w=\frac{1}{\tau}\sum_{j=1}^{m}\Delta t_{j}f(x,u,v_{j})=\sum_{j=1}^{m}q_{j}f(x,u,v_{j}),\quad\text{ where }q_{j}=\frac{\Delta t_{j}}{\tau}.$
$q_{j}\geq 0$ and $\sum_{j=1}^{m}q_{j}=1$, so $w$ is a finite weighted average of points in $S$. Hence,
$w\in\text{Co}(S)=\text{Co}f(x,u,\mathcal{V})$ (20)
Case 2: general measurable case.
For a general measurable $v(\cdot)$, we can approximate it by piecewise-constant $v_{n}(\cdot)$ such that
$f(x,u,v_{n}(\cdot)\rightarrow f(x,u,v(\cdot)\quad\text{ in }L^{1}([t_{k},t_{k+1}]$
This is true because $f(x,u,\cdot)$ is bounded and measurable through $v(\cdot)$.
Define
$w_{n}:=\frac{1}{\tau}\int_{t_{k}}^{t_{k+1}}f(x,u,v_{n}(s)\;ds.$
From Case 1, each $w_{n}\in\text{Co}(S)$.
Moreover,
$|w_{n}-w|$ $=\Big|\frac{1}{\tau}\int_{t_{k}}^{t_{k+1}}(f(x,u,v_{n}(s))-f(x,u,v(s))\;ds\Big|$
$\leq\frac{1}{\tau}\int_{t_{k}}^{t_{k+1}}|f(x,u,v_{n}(s))-f(x,u,v(s))|\;ds\rightarrow 0$
So $w_{n}\rightarrow w$
Finally, $S$ is compact because $\mathcal{V}$ is compact and $f$ is continuous. Hence, Co(S)(S) is compact and therefore closed. Thus, the limit of $w_{n}\in\text{Co}(S)$ also belongs to $\text{Co}(S)$:
$w\in\text{Co}(S)=\text{Co}f(x,u,\mathcal{V})$ (21)
Lemma C.3 (Local Quadratic Error).
For any $k\in\{0,\dots,L-1\}$, any $x$, a fixed control $u\in\mathcal{U}$, any control $v\in\bar{\mathcal{V}}(t_{k})$, there exists some $w\in\text{Co}f(x,u,\mathcal{V})$, such that
$|X_{t_{k+1}}^{t_{k},x,u,v}-(x+\tau w)|\leq C_{1}\tau^{2}$ (22)
@@
We have, from the ODE,
$X(t_{k+1})=X_{t_{k+1}}^{t_{k},x,u,v}=x+\int_{t_{k}}^{t_{k+1}}f(X(s),u,v(s))\;ds.$ (23)
From the definition (18) of $w$,
$x+\tau w=x+\int_{t_{k}}^{t_{k+1}}f(x,u,v(s))ds.$ (24)
Subtract (24) from (23)
$X(t_{k+1})-(x+\tau w)$ $=\int_{t_{k}}^{t_{k+1}}\Big(f(X(s),u,v(s))-f(x,u,v(s)\Big)\;ds$
$\Big|X(t_{k+1})-(x+\tau w)\Big|$ $=\Big|\int_{t_{k}}^{t_{k+1}}\Big(f(X(s),u,v(s))-f(x,u,v(s)\Big)\;ds\Big|$
$\leq\int_{t_{k}}^{t_{k+1}}\Big|f(X(s),u,v(s))-f(x,u,v(s)\Big|\;ds$
From Lipschitz continuity of $f$, $|f(X(s),u,v(s)-f(x,u,v(s))|\leq L_{f}|X(s)-x|$,
$\Big|X(t_{k+1})-(x+\tau w)\Big|\leq\int_{t_{k}}^{t_{k+1}}L_{f}|X(s)-x|\;ds.$
From boundedness of $f$: Let $f||_{\infty}\leq F$. Then for $s\in[t_{k},t_{k+1}]$
$\Big|X(s)-x\Big|=\Big|\int_{t_{k}}^{s}f(X(r),u,v(r)\;dr\Big|\leq\int_{t_{k}}^{s}|f(\cdot)|\;dr\leq\int_{t_{k}}^{s}Fdr=F(s-t_{k}),$
which leads to
$\Big|X(t_{k+1})-(x+\tau w)\Big|$ $\leq\int_{t_{k}}^{t_{k+1}}L_{f}\cdot F\cdot(s-t_{k})\;ds=L_{f}\cdot F\int_{0}^{\tau}r\;dr=\frac{L_{f}\cdot F}{2}\tau^{2}$
which yields equation (22) with $C_{1}=\frac{L_{f}\cdot F}{2}$. ∎
Lemma C.4 (Grid Error).
There exists some constant $C>0$ independent of $\tau$ such that, for any $k\in\{0,\cdots,L\}$, $x\in\mathbb{R}^{d_{x}}$, and $p\in\Delta(I)$, we have
$\max_{\zeta\in\mathcal{Z}}J(t_{k},x,p;\{\eta_{i,\tau}^{\dagger}\},\zeta)\leq V_{\tau}(t_{k},x,p)+C(T-t_{k})\tau$ (25)